# Universal Vessel Segmentation for Multi-Modality Retinal Images

### v3

@@ III-A Overview
We discuss the overall training and inference pipeline (Fig. 2) in this sub-section and provide more details about the major components in the pipeline in the following sub-sections. Given multi-modality retinal images, we first adopt a simple yet widely-used pre-processing method that takes only the green channel from the RGB image. All of the training and inference are then performed on the green-channel images.
@@
Following pre-processing, we train a domain adaptation model for multi-modality retinal images. Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using the pseudo-label methods [42, 76], teacher-student network [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images. Moreover, these methods still require unsupervised finetuning on the target domain images, which deviate from our goal of universal segmentation. Therefore, in this work we propose to use image translation as the preliminary domain adaptation method, which allows us to train a single model for all the modalities.
@@
The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as one type of image. More information of the training dataset is provided in section IV-A1. Furthermore, the translation network is trained with an identity loss (will be discussed in detail in section III-B) which allows us to use a single translation network to transfer all the images to the Topcon CF domain without the need to specify whether they are CF or non-CF. In addition, the translation model also translates CF images from other cameras to Topcon.
@@
After the image translation model is trained, we use it to translate our segmentation training set (19 images from the DRIVE [10] training set with vessel labels) to the Topcon CF domain. The translated images are combined with the original images as an augmented training set to train the downstream segmentation network. Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig. 13). To solve this problem, we propose topology-aware segmentation as a novel domain adaptation method. We show that by learning additional topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains. In practice, in addition to a conventional pixel-wise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
@@
At inference, we can disregard the modality of the image. Given a retinal image from an arbitrary modality, we simply need to (1) extract the green channel; (2) apply the translation network to the green-channel image; (3) apply the topology-aware segmentation network to the translated image to obtain the final segmentation.
@@ III-B Image Translation Network
Multiple GAN-based or Diffusion model-based image translation methods are explored for our task. For Diffusion-based method, we experiment with StyleDiffusion [61] and Unpaired Neuron Schrodinger Bridge (UNSB) [56]. For GAN-based method, we experiment with CycleGAN [44]. We will show in section IV-E2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections discuss the implementation based on CycleGAN and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
@@
To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B. At each iteration, a pair of image A batch and image B batch is selected from the training set. As illustrated by Fig. 3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF) and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain. Both generator A and B are learned with adversarial learning [79, 80], where the generator and discriminator are alternatively updated. At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
@@
where the adversarial loss defined as:
@@ III-C Topology-aware Segmentation Network
The downstream segmentation network is learned with a novel combination of Perceptual Topological Loss $\mathcal{L}_{tc}$ and a Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below accordingly.
@@ III-C1 Perceptual Topological Loss
We first propose to use a perceptual topological loss [62] where the topological information is approximated by the perceptual features encoded by an ImageNet [63] pretrained VGG-16 [30] network. As illustrated by Fig. 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network. In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3 and relu4_3relu4\_3 layers. Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
@@
Figure 4: Illustration for the perceptual topological loss function.
@@ III-C2 Persistent Topological Loss
In 2D digital images, the topology is described by their 0- and 1-homology features, i.e., the connected components and the loops. The persistent homology is a method to describe a process where homology features are created and destroyed in a sequence (called ‘filtration’) of cubical complex [81] (which encodes an image and the higher-dimensional relationships between its pixels to structured cells) and its sub-complexes determined by the values of cells. Although it deviates from the definition and how persistent homology is computed, a simpler way to interpret persistent homology in 2D digital images is by thresholding the image from its highest pixel value to its lowest pixel value and stop at every pixel value that appears in the image. Each thresholded image is a binary image where we can determine the connected components and the loops inside. As demonstrated by the example in Fig. 5, when the threshold value decreases, new pixels are added to the binary image, leading to creation of new homology features or destruction of old features. A homology feature that ‘lives’ through the decreasing thresholds is then called a ‘persistent feature’. We refer the readers to [82] for more details and strict definitions.
Figure 5: A toy example visualizing persistent homology in images: (a) a gray-scale image; (b) the filtration visualized by thresholded binary images. Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinate is (1-Creation Value, 1-Destruction Value).
We propose to use the Spatial-Aware Topological Loss Function (SATLoss) [72] as the persistent topological loss function. After computing persistent diagrams $\mathcal{D}(I_{pred})$ and $\mathcal{D}(I_{GT})$, the persistent features are matched using a spatially-weighted Wasserstein distance [73]:
@@
where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features. $p$ and $\eta(p)$ are a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched with. $q$ is the norm and in practice we take $q=2$. After finding the optimal matching by solving the optimal transport problem, the persistent topological loss function is computed as (reformulating $p$ and $\eta(p)$ by their creation and destruction values):
@@
The afore-hand matching process is based on all the candidate matchings, including matching with points in $\mathcal{D}(I_{GT})$ and matching with the diagonal.
@@ III-C3 Loss Formulation
Finally, the persistent topological loss and the perceptual topological loss are used with a conventional pixel-wise BCELoss:
$\mathcal{L}_{bce}=\frac{1}{n}\sum_{i=1}^{n}y_{i}log(x_{i})+(1-y_{i})log(1-x_{i})$ (10)
where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood value and the ground truth value. $n$ is the total number of pixels in the image. The topological loss functions are controlled by weights $\lambda_{tc}$, $\lambda_{ts}$:
@@ III-C4 Segmentation Network
We use a basic U-Net [31] structure with reduced model size for the segmentation network. In addition, as shown in Fig. 6, we adopt a ‘predict-refine’ strategy which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it. This method has been shown useful in both topology-aware image segmentation [62] and multiple works in deep-learning-based retinal vessel segmentation [41, 42, 43]. In our method, the prediction and refinement networks share the same weight. Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
@@ IV Experiments
Figure 8: Qualitative comparison between our UVSM and finetuned SOTA models on multi-modality retinal images (better zoom in to view). Red, blue orange and green box include the original image, the ground truth label, finetuned FA-Net segmentation and the segmentation from our UVSM. From left to right, up to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR. Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline finetuned models).
TABLE II: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Non - Color Fundus modality (excluding OCTA). The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
Our UVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
TABLE III: Comparison between our UVSM with SOTA finetuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality. The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
Our UVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
TABLE IV: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
@@
Our UVSM 65.40 49.00 65.69 74.32
TABLE V: Comparison between UVSM and general-purpose foundational segmentation models, reporting metric: Dice score.
@@ IV-A1 Image Translation Training
We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k images (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from Heidelberg camera. For Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly-used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos image for FA, MC, FAF modalities, respectively.
@@ IV-A2 Segmentation Training
We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset. In practice, we discard image #14 due to the unusual vessel pattern. In addition, for all the images we resize them to $\times768$ for training. We also resized the ground truth annotation and manually corrected the error caused by interpolation. In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground. In our revised ground truth, we correct these errors and allow more efficient topological learning.
@@ IV-A3 Evaluation
Ten diverse datasets are used for evaluation. Their details can be found in Table I. For the CF modality, we use the public STARE [83], ChaseDB1 [84] and HRF [19] datasets. For MC modality, we use the public IOSTAR [85] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF and JRCIR. Each dataset has 40 $\times768$ images, which is more than most of the commonly-used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF datasets both consist 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retina experts in the UCSD Jacobs Retina Center (JRC). Furthermore, we use an OCTA500-3M and an OCTA600-6M datasets [86] to evaluate the performance of the UVSM on OCTA in two different field-of-views (3mm and 6mm). For the special UWF image (technically UWF is not a modality but only refer to the field-of-view), we use a PRIME-FP20 (MC) dataset [87]. In addition to the ten vessel segmentation dataset, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images on the CF-FA registration dataset [88]. We provide preliminary insights in this paper (Fig. 1, Fig. 8) yet full evaluation requires the vessel labels (unavailable) and is left as future work.

### v4

@@ III-A Overview
We discuss the overall training and inference pipeline (Fig. 2) in this subsection and provide more details about the major components in the pipeline in the following sub-sections. Given multi-modality retinal images, we first adopt a simple yet widely used preprocessing method that takes only the green channel from the RGB image. All of the training and inference are then performed on the green-channel images.
@@
Following preprocessing, we train a domain adaptation model for multi-modality retinal images. Although several domain adaptation methods were proposed for retinal vessel segmentation, such as using pseudo-label methods [42, 76], teacher-student networks [77], adversarial learning [36, 76] and asymmetrical maximum classifier discrepancy [78], they were proposed only to address the domain gap between different datasets (from different cameras) in CF images. Moreover, these methods still require unsupervised fine-tuning on the target domain images, which deviate from our goal of universal segmentation. Therefore, in this work we propose to use image translation as the preliminary domain adaptation method, which allows us to train a single model for all the modalities.
@@
The image translation model aims to translate images from all the modalities to a uniform Topcon CF modality, and is trained between an equivalent number of non-CF images (FA, FAF, MC, IR/NIR) and the Topcon CF images, i.e., during the training, we treat all the non-CF modalities as a single image domain. More information on the training dataset is provided in section IV-A1. Furthermore, the translation network is trained with an identity loss (will be discussed in detail in section III-B), which allows us to use a single translation network to transfer all the images to the Topcon CF domain without the need to specify whether they are CF or non-CF. In addition, the translation model also translates CF images from other cameras to Topcon.
@@
After the image translation model is trained, we use it to translate our segmentation training set (19 images from the DRIVE [10] training set with vessel labels) to the Topcon CF domain. The translated images are combined with the original images as an augmented training set to train the downstream segmentation network. Nevertheless, since we are training a difficult many-to-one (FA, FAF, MC, IR/NIR to CF) translation, the translation model cannot perfectly address the domain gap between the non-CF images and the CF images (Fig. 13). To solve this problem, we propose topology-aware segmentation as a novel domain adaptation method. We show that by learning additional topological features of the vessels, the segmentation model becomes much more robust to inputs from different domains. In practice, in addition to a conventional pixelwise loss function (BCELoss), the segmentation network is jointly trained with two topological loss functions.
@@
At inference, we can disregard the modality of the image. Given a retinal image from an arbitrary modality, we simply need to (1) extract the green channel; (2) apply the translation network to the green channel image; (3) apply the topology-aware segmentation network to the translated image to obtain the final segmentation.
@@ III-B Image Translation Network
Multiple GAN-based and diffusion model-based image translation methods are explored for our task. For the diffusion-based method, we experiment with StyleDiffusion [61] and Unpaired Neuron Schrödinger Bridge (UNSB) [56]. For the GAN-based method, we experiment with CycleGAN [44]. We will show in section IV-E2 that CycleGAN, despite being a classic method, outperforms the other two most recent methods in most datasets. Therefore, we pick CycleGAN as the backbone of our image translation model. Consequently, the following sections discuss the implementation based on CycleGAN, and we refer the readers to [61, 56] for details about the other two methods and to section IV-E2 for our implementation details.
@@
To train the translation network, we treat all the non-CF images (FA, FAF, MC, IR/NIR) as image type A and the Topcon CF images as image type B. At each iteration, one batch of image type A and one batch of image type B are selected from the training set. As illustrated by Fig. 3, generator A is trained to translate all the type A images (non-CF) to a fake type B (CF), and generator B is used to translate type B to a fake type A, which is a neutral non-CF domain. Both generators A and B are learned with adversarial learning [79, 80], where the generator and discriminator are alternately updated. At each iteration, given the image batch $I_{A}$, the generators (A+B) are first updated with the objectives (discriminators frozen):
@@
where the adversarial loss is defined as:
@@ III-C Topology-aware Segmentation Network
The downstream segmentation network is learned with a novel combination of the Perceptual Topological Loss $\mathcal{L}_{tc}$ and the Persistent Topological Loss $\mathcal{L}_{ts}$, introduced below.
@@ III-C1 Perceptual Topological Loss
We first propose to use a perceptual topological loss [62] where the topological information is approximated by the perceptual features encoded by an ImageNet [63]-pretrained VGG-16 [30] network. As illustrated by Fig. 4, image embeddings $\mathcal{E}_{l}(\cdot)$ are computed from the predicted vessel probability map $I_{pred}$ and the ground truth $I_{GT}$, where $l$ denotes different layers in the VGG-16 network. In practice, we take the output of the relu2_2relu2\_2, relu3_3relu3\_3, and relu4_3relu4\_3 layers. Then, the perceptual topological loss function $\mathcal{L}_{tc}$ is defined as:
@@
Figure 4: Illustration of the perceptual topological loss function.
@@ III-C2 Persistent Topological Loss
In a 2D digital image, the topology is described by its 0- and 1-homology features, i.e., the connected components and the loops. The persistent homology is a method to describe a process where homology features are created and destroyed in a sequence (called ‘filtration’) of a cubical complex [81] (which encodes an image and the higher-dimensional relationships between its pixels as structured cells) and its subcomplexes determined by the values of cells. Although it deviates from the formal definition and how persistent homology is computed, a simpler way to interpret persistent homology in 2D digital images is by thresholding the image from its highest pixel value to its lowest pixel value and stopping at every pixel value that appears in the image. Each thresholded image is a binary image, where we can determine the connected components and the loops inside. As demonstrated by the example in Fig. 5, when the threshold value decreases, new pixels are added to the binary image, leading to the creation of new homology features or destruction of old features. A homology feature that ‘lives’ through the decreasing thresholds is then called a ‘persistent feature’. We refer readers to [82] for more details and strict definitions.
Figure 5: A toy example visualizing persistent homology in images: (a) a grayscale image; (b) the filtration visualized by thresholded binary images. Under each image, the digits represent the filtration value (threshold value), the number of 0-homology features, and the number of 1-homology features; (c) the 0-persistent diagram; (d) the 1-persistent diagram.
The persistent homology of an image can be described by a persistent diagram, where each point on the diagram represents one persistent feature, whose coordinates are (1-Creation Value, 1-Destruction Value).
We propose to use a Spatial-Aware Topological Loss Function (SATLoss) [72] as the persistent topological loss function. After computing persistent diagrams $\mathcal{D}(I_{pred})$ and $\mathcal{D}(I_{GT})$, the persistent features are matched using a spatially-weighted Wasserstein distance [73]:
@@
where $s_{p}=\left\|c_{b}(p)-c_{b}(\eta(p))\right\|^{q}$ is a spatial weight that is computed between the creation cells ($c_{b}(\cdot)$) of the persistent features. $p$ and $\eta(p)$ denote a persistent feature from $I_{pred}$ and a candidate persistent feature from $I_{GT}$ to be matched. $q$ is the norm and in practice we take $q=2$. After finding the optimal matching by solving the optimal transport problem, the persistent topological loss function is computed as (reformulating $p$ and $\eta(p)$ by their creation and destruction values):
@@
The aforementioned matching process is based on all the candidate matchings, including matching with points in $\mathcal{D}(I_{GT})$ and matching with the diagonal.
@@ III-C3 Loss Formulation
Finally, the persistent topological loss and the perceptual topological loss are used with a conventional pixelwise BCELoss:
$\mathcal{L}_{bce}=-\frac{1}{n}\sum_{i=1}^{n}y_{i}log(x_{i})+(1-y_{i})log(1-x_{i})$ (10)
where $x_{i}$ and $y_{i}$ are the per-pixel predicted likelihood values and the ground truth value. $n$ is the total number of pixels in the image. The topological loss functions are controlled by weights $\lambda_{tc}$ and $\lambda_{ts}$:
@@ III-C4 Segmentation Network
We use a basic U-Net [31] structure with reduced model size for the segmentation network. In addition, as shown in Fig. 6, we adopt a ‘predict-refine’ strategy, which first makes a coarse prediction of the vessels and feeds the initial prediction and the input image to the network again to refine it. This method has been shown to be useful in both topology-aware image segmentation [62] and multiple works in deep learning-based retinal vessel segmentation [41, 42, 43]. In our method, the prediction and refinement networks share the same weights. Additionally, at both the initial prediction stage and the refinement stage, the loss functions are computed to optimize the network.
@@ IV Experiments
TABLE II: Training (one-time) and inference (per item, image size 768 $\times$ 768) time of the URVSM: translation (Trans.) + segmentation (Seg.). Device: i9-13900K CPU + 1 RTX 4090 GPU.
Trans. Train Seg. Train Trans. Inference Seg. Inference
6.9h 11.0h 0.15s 0.07s
Figure 8: Qualitative comparison between our URVSM and fine-tuned SOTA models on multi-modality retinal images (best view with zooming in). Red, blue, orange and green box include the original image, the ground truth label, fine-tuned FA-Net segmentation and the segmentation from our URVSM. From left to right, top to down: Topcon CF, Canon CF, Zeiss CF, Heidelberg MC, Optos FA, Heidelberg FA, Optos FAF, Heidelberg FAF, Heidelberg IR. Optovue OCTA 3mm, Optovue OCTA 6mm, Bottom right shows the qualitative results for Optos MC and Canon FA (no ground truth available to compare with and to train the baseline fine-tuned models).
TABLE III: Comparison between our URVSM and SOTA fine-tuned (unsupervised domain adaptation and supervised) methods on Non Color Fundus modality (excluding OCTA). The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
Our URVSM 96.10 74.66 97.84 75.31 72.85 73.66 71.81 76.30 28.8 5.5
TABLE IV: Comparison between our URVSM with SOTA fine-tuned (unsupervised domain adaptation and supervised) methods on Color Fundus modality. The green boxes indicate the best result and the blue boxes indicate the second-best result.
@@
Our URVSM 96.55 77.34 98.05 77.81 75.38 76.34 74.63 78.34 70.1 7.5
TABLE V: Performance on Ultra-Wide-Field retinal images, pure zero-shot performance on OCTA, and comparison with SOTA supervised methods.
@@
Our URVSM 65.40 49.00 65.69 74.32
TABLE VI: Comparison between URVSM and general-purpose foundational segmentation models, metric reported: Dice score.
@@ IV-A1 Image Translation Training
We prepared a multi-modality database with over 150k de-identified images from the database of the Jacobs Retina Center at UCSD. However, in our previous numerous attempts we found it very difficult and time-consuming to tune the training of the image translation network (regardless of the methods) for a large dataset. Therefore, in this work, we use a 4k-image (2k CF, 500 FA, FAF, MC, IR/NIR, respectively) dataset from the larger 150k database. Each image has a resolution of 768 $\times$ 768. All the CF images are from the 30-degree (field-of-view) Topcon camera and all the IR/NIR images are from the Heidelberg camera. For FA, FAF and MC, 250 images are from the Heidelberg camera. For the Optos camera, we crop random regions from the original 135 degrees to get 30-40-degree conventional field-of-view images (so that the image size is the same as the other modalities) and apply a commonly used circular mask (Fig. 7(a)). In the dataset, there are 250 cropped and masked Optos images for FA, MC, FAF modalities, respectively.
@@ IV-A2 Segmentation Training
We train the segmentation network using only the 20 training images from the public DRIVE [10] CF dataset. In practice, we discard image #14 due to the unusual vessel pattern. In addition, for all the images we resize them to $\times768$ for training. We also resized the ground truth annotations and manually corrected the errors caused by interpolation. In addition, according to our examination, the original DRIVE ground truth has multiple topological errors such as small redundant holes in the vessel foreground. In our revised ground truth, we correct these errors and allow more efficient topological learning.
@@ IV-A3 Evaluation
Ten diverse datasets were used for evaluation. Their details can be found in Table I. For the CF modality, we use the public STARE [83], ChaseDB1 [84], and HRF [19] datasets. For MC modality, we use the public IOSTAR [85] dataset. For FA, FAF and IR/NIR, we prepared three new datasets: JRCFA, JRCFAF, and JRCIR. Each dataset has 40 $\times768$ images, which is more than most of the commonly used public retinal vessel segmentation (CF) datasets. The JRCFA and JRCFAF datasets each consist of 20 images from Heidelberg and 20 from Optos, respectively. The JRCIR dataset consists of 40 Heidelberg (the only camera for this modality) images. All the vessel annotations are either labeled or carefully examined and refined by the retinal experts in the UCSD Jacobs Retina Center (JRC). Furthermore, we use an OCTA500-3M and an OCTA600-6M dataset [86] to evaluate the performance of the URVSM on OCTA in two different fields of view (3mm and 6mm). For the special UWF images (technically UWF refers to field-of-view rather than a distinct modality), we use the PRIME-FP20 (MC) dataset [87]. In addition to the ten vessel segmentation datasets, we also evaluate our model qualitatively on the Optos MC images using images from our 150K database and on the Canon FA images from the CF-FA registration dataset [88]. We provide preliminary insights in this paper (Fig. 1, Fig. 8), yet full evaluation requires the vessel labels (unavailable) and is left as future work. Additionally, we use the Massachusetts Roads [89] (road segmentation in aerial imagery) and CREMI [90] (extracellular boundary segmentation in electronic microscopy) datasets for the extended experiments on topological domain adaptation on more general curvilinear structure segmentation. The test set of Massachusetts Roads and all slices in volume C in CREMI were used for evaluation.